OpenStudy (anonymous):

olve for x, given the equation Square root of x plus 9 − 4 = 1. im so confused on this topic can someone help me?

OpenStudy (anonymous):

\[\sqrt{x+9-4=1}\]

OpenStudy (jtvatsim):

\[\sqrt{x+9} -4 = 1\] I think you mean this?